# Geometric inequalities and the Alexandrov-Bakelman-Pucci technique

### 12 Mar 2026 v1

@@ 1. The classical isoperimetric inequality and the Sobolev inequality
In [11],[12], Cabré gave an elegant proof of the isoperimetric inequality based on what has become known as the Alexandrov-Bakelman-Pucci technique. The Alexandrov-Bakelman-Pucci technique in Cabré’s work is, in a certain sense, dual to the optimal transport approach developed by McCann and Trudinger. In the optimal transport approach, one considers a convex function satisfying $\det D^{2}u=1$ and concludes that $\Delta u\geq n$ by the arithmetic-geometric mean inequality. In the Alexandrov-Bakelman-Pucci approach, one considers a solution of $\Delta u=n$ and concludes that $\det D^{2}u\leq 1$ at each point where the Hessian of $u$ is weakly positive definite. The Alexandrov-Bakelman-Pucci technique only requires standard existence theorems for linear elliptic PDE.
@@
for each point $x\in\partial D$. Consequently, we can find a point $\bar{x}\in D\setminus\partial D$ where the function $w$ attains its minimum. Clearly, $\nabla w(\bar{x})=0$ and $D^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi=\nabla u(\bar{x})$ and $D^{2}u(\bar{x})\geq 0$. This implies $\bar{x}\in A$ and $\Phi(\bar{x})=\xi$. Thus, $B^{n}\subset\Phi(A)$.
@@ 2. The Fenchel-Willmore-Chen inequality for submanifolds of Euclidean space
Proof. Let us fix an arbitrary vector $\xi\in\mathbb{R}^{n+m}$ such that $|\xi|<1$. We define a function $w:\Sigma\to\mathbb{R}$ by $w(x):=-\langle x,\xi\rangle$. We can find a point $\bar{x}\in\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi$ is orthogonal to the tangent space $T_{\bar{x}}\Sigma$ and $\langle I\!I(\bar{x}),\xi\rangle\geq 0$. Therefore, if we define a normal vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ by $\bar{y}=\xi$, then $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$.
@@
The assertion follows by taking the determinant.
@@ 3. The Sobolev inequality for submanifolds of Euclidean space
In this section, we explain how the Alexandrov-Bakelman-Pucci technique can be used to prove a isoperimetric inequality for submanifolds.
@@
for each point $x\in\partial\Sigma$. Consequently, we can find a point $\bar{x}\in\Sigma\setminus\partial\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. Therefore, we can find a vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ such that $\xi=\nabla^{\Sigma}u(\bar{x})+\bar{y}$ and $D_{\Sigma}^{2}u(\bar{x})-\langle I\!I(\bar{x}),\bar{y}\rangle\geq 0$. This implies $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$.
@@
The assertion follows by taking the determinant.
@@ 5. The Sobolev inequality for manifolds with nonnegative Ricci curvature
Using (12) and the inequality $\sup_{x\in D}d(x,p)<r$, we conclude that the point $\bar{x}$ lies in the interior of $D$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. The path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=p$. The first variation formula implies that
@@
Since $d(\bar{x},p)<r$, it follows that $|\nabla u(\bar{x})|<1$. Therefore, $\bar{x}\in U$. Finally, for each point $x\in D$, we have
@@
$=r\,u(\bar{x})+\frac{1}{2}\,r^{2}\,|\nabla u(\bar{x})|^{2}.$
Thus, $\bar{x}\in A_{r}$ and $\Phi_{r}(\bar{x})=p$. This completes the proof of Lemma 5.4.
@@
Proof. Since $\bar{x}\in A_{r}$, the path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula.
@@
This contradicts Lemma 5.5.
@@
Finally, the differential of $\Phi_{t}$ at the point $\bar{x}$ maps the vector $e_{i}\in T_{\bar{x}}M$ to the vector $X_{i}(t)\in T_{\bar{\gamma}(t)}M$. Consequently, $|\det D\Phi_{t}(\bar{x})|=\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows.
@@
Proof. Since $\lim_{t\to 0}|\det D\Phi_{t}(x)|=1$, the assertion follows from Proposition 5.7.
@@ 6. The Fenchel-Willmore-Chen inequality for hypersurfaces in manifolds with nonnegative Ricci curvature
Then $d(\bar{x},p)<r$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. The path $\bar{\gamma}$ minimizes the functional $\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=p$. The first variation formula implies that $\bar{\gamma}^{\prime}(0)\in T_{\bar{x}}^{\perp}\Sigma$. We define $\bar{y}=\bar{\gamma}^{\prime}(0)$. Then
@@
Since $d(\bar{x},p)<r$, it follows that $|\bar{y}|<1$. Therefore, $(\bar{x},\bar{y})\in U$. Finally, for each point $x\in\Sigma$, we have
$d\big(x,\exp_{\bar{x}}(r\bar{y})\big)=d(x,p)\geq d(\bar{x},p)=r\,|\bar{y}|.$
Thus, $(\bar{x},\bar{y})\in A_{r}$ and $\Phi_{r}(\bar{x},\bar{y})=p$. This completes the proof of Lemma 6.2.
@@
Proof. Since $(\bar{x},\bar{y})\in A_{r}$, the path $\bar{\gamma}$ minimizes the functional $\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula.
@@
This contradicts Lemma 5.5.
@@
Finally, it is easy to see that $|\det D\Phi_{t}(\bar{x},\bar{y})|=t\,\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows.
@@
Proof. Since $\lim_{t\to 0}t^{-1}\,|\det D\Phi_{t}(x,y)|=1$, the assertion follows from Proposition 6.5.

### 18 Mar 2026 v2

@@ 1. The classical isoperimetric inequality and the Sobolev inequality
In [11],[12], Cabré gave an elegant proof of the isoperimetric inequality based on what has become known as the Alexandrov-Bakelman-Pucci technique. The Alexandrov-Bakelman-Pucci technique in Cabré’s work is, in a certain sense, dual to the optimal transport approach developed by McCann and Trudinger. In the optimal transport approach, one considers a convex function satisfying $\det D^{2}u=1$ and concludes that $\Delta u\geq n$ by the arithmetic-geometric mean inequality. In the Alexandrov-Bakelman-Pucci approach, one considers a solution of $\Delta u=n$ and concludes that $\det D^{2}u\leq 1$ at each point where the Hessian of $u$ is weakly positive definite. The Alexandrov-Bakelman-Pucci technique only requires existence and regularity results for linear partial differential equations.
@@
for each point $x\in\partial D$. Consequently, we can find a point $\bar{x}\in D\setminus\partial D$ where the function $w$ attains its minimum. Clearly, $\nabla w(\bar{x})=0$ and $D^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi=\nabla u(\bar{x})$ and $D^{2}u(\bar{x})\geq 0$. This implies $\bar{x}\in A$ and $\Phi(\bar{x})=\xi$. Thus, $B^{n}\subset\Phi(A)$. This completes the proof of Lemma 1.3.
@@ 2. The Fenchel-Willmore-Chen inequality for submanifolds of Euclidean space
Proof. Let us fix an arbitrary vector $\xi\in\mathbb{R}^{n+m}$ such that $|\xi|<1$. We define a function $w:\Sigma\to\mathbb{R}$ by $w(x):=-\langle x,\xi\rangle$. We can find a point $\bar{x}\in\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. From this, we deduce that $\xi$ is orthogonal to the tangent space $T_{\bar{x}}\Sigma$ and $\langle I\!I(\bar{x}),\xi\rangle\geq 0$. Therefore, if we define a normal vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ by $\bar{y}=\xi$, then $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$. This completes the proof of Lemma 2.2.
@@
The assertion follows by taking the determinant. This completes the proof of Lemma 2.3.
@@ 3. The Sobolev inequality for submanifolds of Euclidean space
In this section, we explain how the Alexandrov-Bakelman-Pucci technique can be used to prove an isoperimetric inequality for submanifolds.
@@
for each point $x\in\partial\Sigma$. Consequently, we can find a point $\bar{x}\in\Sigma\setminus\partial\Sigma$ where the function $w$ attains its minimum. Clearly, $\nabla^{\Sigma}w(\bar{x})=0$ and $D_{\Sigma}^{2}w(\bar{x})\geq 0$. Therefore, we can find a vector $\bar{y}\in T_{\bar{x}}^{\perp}\Sigma$ such that $\xi=\nabla^{\Sigma}u(\bar{x})+\bar{y}$ and $D_{\Sigma}^{2}u(\bar{x})-\langle I\!I(\bar{x}),\bar{y}\rangle\geq 0$. This implies $(\bar{x},\bar{y})\in A$ and $\Phi(\bar{x},\bar{y})=\xi$. Thus, $B^{n+m}\subset\Phi(A)$. This completes the proof of Lemma 3.4.
@@
The assertion follows by taking the determinant. This completes the proof of Lemma 3.5.
@@ 5. The Sobolev inequality for manifolds with nonnegative Ricci curvature
Using (12) and the inequality $\sup_{x\in D}d(x,p)<r$, we conclude that the point $\bar{x}$ lies in the interior of $D$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. Clearly, $r\,|\bar{\gamma}^{\prime}(t)|=d(\bar{x},p)$ for all $t\in[0,r]$. Moreover,
$r\,u(\gamma(0))+\frac{1}{2}\,r\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ $\geq r\,u(\gamma(0))+\frac{1}{2}\,d(\gamma(0),p)^{2}$
$\geq r\,u(\bar{x})+\frac{1}{2}\,d(\bar{x},p)^{2}$
for every smooth path $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=p$. Therefore, the path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=p$. The first variation formula implies that
@@
Since $d(\bar{x},p)<r$, it follows that $|\nabla u(\bar{x})|<1$. Therefore, $\bar{x}\in U$. Finally,
@@
$=r\,u(\bar{x})+\frac{1}{2}\,r^{2}\,|\nabla u(\bar{x})|^{2}$
for each point $x\in D$. Thus, $\bar{x}\in A_{r}$ and $\Phi_{r}(\bar{x})=p$. This completes the proof of Lemma 5.4.
@@
Proof. Since $\bar{x}\in A_{r}$, we obtain
$r\,u(\gamma(0))+\frac{1}{2}\,r\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ $\geq r\,u(\gamma(0))+\frac{1}{2}\,d(\gamma(0),\bar{\gamma}(r))^{2}$
$\geq r\,u(\bar{x})+\frac{1}{2}\,r^{2}\,|\nabla u(\bar{x})|^{2}$
for every smooth path $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=\bar{\gamma}(r)$. Therefore, the path $\bar{\gamma}$ minimizes the functional $u(\gamma(0))+\frac{1}{2}\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in D$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula. This completes the proof of Lemma 5.5.
@@
This contradicts Lemma 5.5. This completes the proof of Lemma 5.6.
@@
Finally, the differential of $\Phi_{t}$ at the point $\bar{x}$ maps the vector $e_{i}\in T_{\bar{x}}M$ to the vector $X_{i}(t)\in T_{\bar{\gamma}(t)}M$. Consequently, $|\det D\Phi_{t}(\bar{x})|=\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows. This completes the proof of Proposition 5.7.
@@
Proof. Since $\lim_{t\to 0}|\det D\Phi_{t}(x)|=1$, the assertion follows from Proposition 5.7. This completes the proof of Corollary 5.8.
@@ 6. The Fenchel-Willmore-Chen inequality for hypersurfaces in manifolds with nonnegative Ricci curvature
Then $d(\bar{x},p)<r$. Let $\bar{\gamma}:[0,r]\to M$ be a minimizing geodesic such that $\bar{\gamma}(0)=\bar{x}$ and $\bar{\gamma}(r)=p$. Clearly, $r\,|\bar{\gamma}^{\prime}(t)|=d(\bar{x},p)$ for all $t\in[0,r]$. Moreover,
$r\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt\geq d(\gamma(0),p)^{2}\geq d(\bar{x},p)^{2}$
for every smooth path $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=p$. Therefore, the path $\bar{\gamma}$ minimizes the functional $\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=p$. The first variation formula implies that $\bar{\gamma}^{\prime}(0)\in T_{\bar{x}}^{\perp}\Sigma$. We define $\bar{y}=\bar{\gamma}^{\prime}(0)$. Then
@@
Since $d(\bar{x},p)<r$, it follows that $|\bar{y}|<1$. Therefore, $(\bar{x},\bar{y})\in U$. Finally,
$d\big(x,\exp_{\bar{x}}(r\bar{y})\big)=d(x,p)\geq d(\bar{x},p)=r\,|\bar{y}|$
for each point $x\in\Sigma$. Thus, $(\bar{x},\bar{y})\in A_{r}$ and $\Phi_{r}(\bar{x},\bar{y})=p$. This completes the proof of Lemma 6.2.
@@
Proof. Since $(\bar{x},\bar{y})\in A_{r}$, we obtain
$r\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt\geq d(\gamma(0),\bar{\gamma}(r))^{2}\geq r^{2}\,|\bar{y}|^{2}$
for every smooth path $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=\bar{\gamma}(r)$. Therefore, the path $\bar{\gamma}$ minimizes the functional $\int_{0}^{r}|\gamma^{\prime}(t)|^{2}\,dt$ among all smooth paths $\gamma:[0,r]\to M$ satisfying $\gamma(0)\in\Sigma$ and $\gamma(r)=\bar{\gamma}(r)$. Hence, the assertion follows from the second variation formula. This completes the proof of Lemma 6.3.
@@
This contradicts Lemma 6.3. This completes the proof of Lemma 6.4.
@@
Finally, it is easy to see that $|\det D\Phi_{t}(\bar{x},\bar{y})|=t\,\det P(t)$ for all $t\in(0,r)$. Putting these facts together, the assertion follows. This completes the proof of Proposition 6.5.
@@
Proof. Since $\lim_{t\to 0}t^{-1}\,|\det D\Phi_{t}(x,y)|=1$, the assertion follows from Proposition 6.5. This completes the proof of Corollary 6.5.